OpenStudy (anonymous):

What is the inverse of the given relation? y=3x+12

OpenStudy (mimi_x3):

Switch x with y and then solve for y Switch x and y: x=3y+12 Isolate y: x-12=3y \[\frac{x-12}{3}=y\]

OpenStudy (mimi_x3):

@crehgio @DanJS @dan815 @nincompoop @calculusxy @camerondoherty @Nnesha PLease medal me <3

OpenStudy (mimi_x3):

@bibby @Kainui @iambatman

OpenStudy (calculusxy):

Great explanation! @Mimi_x3

OpenStudy (mimi_x3):

Thanks @calculusxy <3

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (dan815):

this is abuse, you are not allowed to ask for medals, i am going to have to report you

OpenStudy (mimi_x3):

ffs @dan815 STOP HARASSING ME OR I WILL CALL MODS

OpenStudy (dan815):

go ahead, it will only help me bcome a mod faster, then ill haunt you